# V2X-M2C: Efficient Multi-Module Collaborative Perception with Two Connections

### v1

@@ 3.5 Head
After passing the last depth, we get the final output $\mathbf{M}^{D}$. We extract the ego agent output $\mathbf{M}^{D}_{1}$ and apply two $1\times 1$ convolution layers, $f^{\mathrm{cls}}_{\mathrm{head}}(\cdot)$ and $f^{\mathrm{reg}}_{\mathrm{head}}(\cdot)$, for the detection box classification and regression, respectively. Specifically, $f^{\mathrm{cls}}_{\mathrm{head}}(\cdot)$ makes the classification output tensor $\mathbf{\hat{Y}}_{\mathrm{cls}}\in\mathbb{R}^{H\times W\times 2}$ that identifies whether an object is a vehicle or not, and $f^{\mathrm{reg}}_{\mathrm{head}}(\cdot)$ makes the regression output tensor $\mathbf{\hat{Y}}_{\mathrm{reg}}\in\mathbb{R}^{H\times W\times 14}$ that contains center of boxes $(x,y,z)$, size of boxes $(w,l,h)$ and heading of vehicle $\phi$. We use the total loss $\mathcal{L}_{\mathrm{total}}$ = $\mathcal{L}_{\mathrm{cls}}+\mathcal{L}_{\mathrm{reg}}$ where $\mathcal{L}_{\mathrm{cls}}$ is calculated using $\mathbf{\hat{Y}}_{\mathrm{cls}}$ and focal loss [27] and $\mathcal{L}_{\mathrm{reg}}$ is calculated using $\mathbf{\hat{Y}}_{\mathrm{reg}}$ and the smooth L1 loss [31].
@@ 4 Experiments
Comparison Models and Datasets. We compare our proposed architecture with state-of-the-art (SOTA) models [1], [2], [3] on the V2XSet [2] and OPV2V [18] datasets to validate performance in both the V2V (Vehicle-to-Vehicle) and V2X scenarios. V2XSet is a simulated dataset, including V2X scenarios, co-simulated using CARLA [19] and OpenCDA [20]. OPV2V is a simulated dataset, only including V2V scenarios. We set the LiDAR detection range as $x\in[-140.8,140.8]$ and $y\in[-38.4,38.4]$. A detailed description of the datasets and models is in Appendix B.1 and B.2.
@@
Evaluation Metrics and Noise Settings. We calculate the Average Precision (AP) at the Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 to evaluate 3D detection accuracy and use a number of parameters (# Params) and GFLOPs to compare model efficiency. To compare models for various situations, we add noise to training and inference data, which includes latency, localization, and heading noises. The latency $t_{\mathrm{lag.}}$ is divided into constant or random, and the random $t_{\mathrm{lag.}}$ is based on the total time delay [2]. The localization and heading noises use normal distributions with standard deviations $\sigma_{\mathrm{loc.}}$ and $\sigma_{\mathrm{hdg.}}$, respectively. We train models with three different noisy settings: perfect, simple noise, and harsh noise. The perfect setting does not apply any noise, and the simple noise setting applies constant $t_{\mathrm{lag.}}$ of 100 ms, and the localization and heading noises with $\sigma_{\mathrm{loc.}}=0.2\,$m and $\sigma_{\mathrm{hdg.}}=$$$ 0.2  °. Detailed information about training and inference setting is described in Appendix B.4 and  B.5.
Table 1: Comparison of detection accuracy and model efficiency.
Model Perfect Simple Noise # Params GFLOPs
V2XSet OPV2V V2XSet OPV2V
AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7
No Fusion 0.754 / 0.602 0.659 / 0.545 0.754 / 0.602 0.659 / 0.545 - -
Late Fusion 0.885 / 0.764 0.852 / 0.768 0.844 / 0.552 0.821 / 0.597 6.58M 25
CoBEVT [1] 0.896 / 0.841 0.855 / 0.793 0.870 / 0.740 0.780 / 0.615 11.14M 213
V2X-ViT [2] 0.886 / 0.807 0.848 / 0.769 0.872 / 0.734 0.795 / 0.632 14.17M 287
Where2comm [3] 0.913 / 0.853 0.854 / 0.789 0.888 / 0.751 0.778 / 0.624 8.69M 177
M2C-S (Ours) 0.922 / 0.860 0.887 / 0.828 0.907 / 0.769 0.844 / 0.682 5.97M 103
M2C-P (Ours) 0.904 / 0.853 0.869 / 0.816 0.892 / 0.779 0.828 / 0.681 5.92M 103
Figure 4: Robustness in various noise ranges.
@@ 4.2 Quantitative evaluation
Detection Accuracy. Table 1 compares the proposed V2X-M2C with the other SOTA methods. In particular, we consider single-agent detection (No Fusion), late fusion, CoBEVT [1], V2X-ViT [2], and Where2comm [3] (see detection visualization in Appendix C.2). We observe that: i) Compared with No Fusion, M2C-S/M2C-P enhances the accuracy by 42.70%/41.62% on V2XSet, and 51.94%/49.75% on OPV2V in the perfect setting. ii) In the perfect setting, M2C-S increases the accuracy by 0.77% compared with Where2comm on V2XSet and 4.49% compared with CoBEVT on OPV2V at AP@0.7. iii) In the simple noise setting, M2C-S/M2C-P increases the accuracy by 2.28%/3.63% compared with Where2comm on V2XSet, and 8.00%/7.79% compared with V2X-ViT on OPV2V at AP@0.7. Our proposed M2C-S and M2C-P outperform the SOTA models on both datasets regardless of noise.
Model Efficiency. Table 1 compares model efficiency in terms of # Params and GFLOPs. The # Params in M2C-S/M2C-P is reduced by 58.1%/58.4% compared with V2X-ViT and by 31.65%/32.22% compared with Where2comm [3]. The GFLOPs of M2C-S and M2C-P are reduced by 64.11% compared with V2X-ViT [2] and 41.80% compared with Where2comm [3]. Substantially, we find that our models have fewer parameters than the PoinPillars-based late fusion. This shows our models can work in lightweight hardware.
Noise Robustness. To evaluate the robustness to noises, we utilize models trained with a harsh noise setting. We apply various ranges of noise components: random $t_{\mathrm{lag.}}\in[0\,\mathrm{ms},500\,\mathrm{ms}]$, $\sigma_{\mathrm{loc.}}\in[0.0\,\mathrm{m},0.5\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg.}}\in[$$,$$]$ 0.0  ° , 1.0  ° ] from the zero noise (0 ms, 0.0 m, 0.0°0.0°0.0  °). As shown in Fig. 4, M2C-S and M2C-P outperform the SOTA models in all the noise components. At the same time, M2C-S and M2C-P show the lowest noise sensitivity among the SOTA models (Detailed values in Appendix C.1). For the severe noise setting (500 ms / 0.5 m / 1.0°1.0°1.0  °), the rate of change (i.e., sensitivity) of M2C-S and M2C-P decreases by 8.82%/4.40%/12.52% and 9.07%/4.16%/12.79% from the zero noise, respectively, while that of V2X-ViT [2] decreases by 11.34%/4.54%/13.04% and that of Where2comm [3] decreases by 9.96%/5.51%/13.71%.
@@ 4.3 Qualitative Evaluation
Visualization of CCL Weights. In Figure 5, we visualize the weights of CCL in M2C-P that consist of four sections indicating A-Att, S-Att, H-conv, and original layers for identifying the tendency of learning. We observe that i) the weights of H-Conv show the highest values in the perfect and simple noise setting as the depth gets deeper. In particular, in the third depth, the sum of the weights of the H-Conv section is 482.498 in the perfect setting and 541.413 in the simple noise setting, while that of the other sections (A-Att / S-Att / Original) is 210.488 / 275.469 / 300.404 and 185.661 / 260.079 / 298.485, respectively. ii) Comparing the perfect to noise setting, the sum of the weights of the H-Conv section increases from 482.498 to 541.413, whereas the other sums decrease. It highlights that H-Conv plays a main role in compensating for the other intents and overcoming noise disturbance. This also contributes to the low noise sensitivity of our models discussed in Section 4.2.
(a) Perfect Setting             (b) Noise Setting
Figure 5: Visualization of CCL weights in M2C-P with the perfect and noisy setting. The weight matrix is divided into four sections (A-Att, S-Att, H-conv, original). The number in the parentheses of each section is the sum of the section’s weights.
(a) V2X Scenario       (b) V2X-ViT      (c) M2C-S        (d) M2C-P

### v2

@@ 3.5 Head
After passing the last depth, we get the final output $\mathbf{M}^{D}$. We extract the ego agent output $\mathbf{M}^{D}_{1}$ and apply two $1\times 1$ convolution layers, $f^{\mathrm{cls}}_{\mathrm{head}}(\cdot)$ and $f^{\mathrm{reg}}_{\mathrm{head}}(\cdot)$, for the detection box classification and regression, respectively. Specifically, $f^{\mathrm{cls}}_{\mathrm{head}}(\cdot)$ makes the classification output tensor $\mathbf{\hat{Y}}_{\mathrm{cls}}\in\mathbb{R}^{H\times W\times 2}$ that identifies whether an object is a vehicle or not, and $f^{\mathrm{reg}}_{\mathrm{head}}(\cdot)$ makes the regression output tensor $\mathbf{\hat{Y}}_{\mathrm{reg}}\in\mathbb{R}^{H\times W\times 14}$ that contains center of boxes $(x,y,z)$, size of boxes $(w,l,h)$ and heading of vehicle $\phi$. We use the total loss $\mathcal{L}_{\mathrm{total}}$ = $\mathcal{L}_{\mathrm{cls}}+\mathcal{L}_{\mathrm{reg}}$ where $\mathcal{L}_{\mathrm{cls}}$ is calculated using $\mathbf{\hat{Y}}_{\mathrm{cls}}$ and focal loss (Lin et al. [2017]) and $\mathcal{L}_{\mathrm{reg}}$ is calculated using $\mathbf{\hat{Y}}_{\mathrm{reg}}$ and the smooth L1 loss (Berrada et al. [2018]).
@@ 4.1 Experimental Settings
Comparison Models and Datasets. We compare our proposed architecture with state-of-the-art (SOTA) models, CoBEVT Xu et al. [2022a], V2X-ViT Xu et al. [2022b], Where2comm Hu et al. [2022], CoAlign Lu et al. [2023], on the datasets of V2XSet (Xu et al. [2022b]), OPV2V (Xu et al. [2022c]), and DAIR-V2X (Yu et al. [2022]). V2XSet is a simulated dataset, including V2X scenarios, using the CARLA simulator (Dosovitskiy et al. [2017]) and OpenCDA (Xu et al. [2021]). OPV2V is a simulated dataset, only including V2V scenarios. Also, DAIR-V2X is a real-world dataset, including V2X scenarios, collected from one actual vehicle and one road infrastructure. We set the LiDAR detection range as $x\in[-140.8,140.8]$ and $y\in[-38.4,38.4]$ on V2XSet and OPV2V, and $x\in[-102.4,102.4]$ and $y\in[-38.4,38.4]$ on DAIR-V2X. A detailed description of the datasets and models is in Appendix A.1 and A.2.
@@
Evaluation Metrics and Noise Settings. We calculate the Average Precision (AP) at the Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 to evaluate 3D detection accuracy and use the number of parameters (# Params) and GFLOPs to compare model efficiency. To consider communication noise while sharing the feature of connected agents, we add communication noise to training and inference data, which includes latency, heading noises, and localization noises. The latency is based on the total time delay used in Xu et al. [2022b] and is set to follow uniform distribution $U(0,t_{\mathrm{lag.}})$ with the maximum latency $t_{\mathrm{lag.}}$. The heading and localization noises follow normal distributions with zero mean and standard deviations $\sigma_{\mathrm{hdg.}}$ and $\sigma_{\mathrm{loc.}}$, that is, $\mathcal{N}(0,\sigma_{\mathrm{hdg.}}^{2})$ and $\mathcal{N}(0,\sigma_{\mathrm{loc.}}^{2})$, respectively. We use three different noise settings: perfect, mild noise, and harsh noise. In the perfect setting, we let $t_{\mathrm{lag.}}=0,\sigma_{\mathrm{hdg.}}=0,\sigma_{\mathrm{loc.}}=0$, and in the mild noise setting, we let $t_{\mathrm{lag.}}=200,\sigma_{\mathrm{hdg.}}=0.2,\sigma_{\mathrm{loc.}}=0.2$. In the harsh noise setting, we use different values of $t_{\mathrm{lag.}}\in[0,500],\sigma_{\mathrm{hdg.}}=[0,1.0],\sigma_{\mathrm{loc.}}=[0,0.5]$. Also, we train three types of models based on the noise settings: perfect, noise, and fine-tuned. The perfect model is trained in the perfect setting, the noise model in the simple noise setting (see Appendix A.3 for details), and the fine-tuned model fine-tunes the perfect model with the mild noise setting. We validate the perfect model, noise model, and fine-tuned model with the perfect setting, mild noise setting, and harsh noise setting, respectively.
Training Parameters. For a fair comparison, we train both our models and SOTA models using AdamW (Loshchilov & Hutter [2017]), with a learning rate of 3e-4 and a weight decay of 0.01. We use the same learning scheduler, Cosine Annealing Warm-Up Restarts (Loshchilov & Hutter [2016]), applying a warm-up learning rate of 2e-4. Regarding V2XSet and OPV2V, we train the models for up to 40 epochs and 10 warm-up epochs. We also train the models for up to 20 epochs and 5 warm-up epochs for DAIR-V2X. The models are trained and validated on an RTX 4090.
Model Parameters. After the feature extraction, we compress the channel size of the feature from $C=256$ to $C/N=8$ for the feature sharing. We set the channel size of multi-agent feature $\mathbf{M^{0}}$ as 256. Regarding the fusion module, HRPE in A-att utilizes the relative angle $\theta$ and distance $d$. For V2XSet and OPV2V, we divide $\theta$ and $d$ into 20°20°20  ° and 25 m, and into 10°10°10  ° and 15 m for DAIR-V2X. In S-Att, the lightweight DiNAT (Hassani & Shi [2022]) features two depths. For DiNAT’s hyperparameters, we use a $7\times 7$ kernel with dilation rates of 4 and 2 applied to each depth, respectively. The S-Conv comprises three blocks, each consisting of a convolution layer with a $3\times 3$ kernel.
Table 1: Comparison of detection accuracy on V2XSet, OPV2V, and DAIR-V2X.
Model Perfect Mild Noise
V2XSet OPV2V DAIR-V2X V2XSet OPV2V DAIR-V2X
AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7
No Fusion 0.754 / 0.602 0.659 / 0.545 0.562 / 0.452 0.754 / 0.602 0.659 / 0.545 0.562 / 0.452
Late Fusion 0.885 / 0.764 0.852 / 0.768 0.573 / 0.397 0.844 / 0.552 0.726 / 0.539 0.549 / 0.364
CoBEVT 0.880 / 0.825 0.852 / 0.787 0.570 / 0.452 0.787 / 0.600 0.698 / 0.522 0.550 / 0.407
V2X-ViT 0.866 / 0.773 0.843 / 0.763 0.603 / 0.455 0.811 / 0.648 0.755 / 0.615 0.572 / 0.409
Where2comm 0.904 / 0.835 0.858 / 0.789 0.594 / 0.429 0.852 / 0.668 0.760 / 0.576 0.562 / 0.384
CoAlign 0.888 / 0.799 0.819 / 0.741 0.578 / 0.433 0.816 / 0.605 0.726 / 0.539 0.557 / 0.397
ParCon (Ours) 0.891 / 0.839 0.888 / 0.830 0.613 / 0.482 0.850 / 0.707 0.790 / 0.649 0.596 / 0.438
Table 2: Comparison of efficiency.
Model # Params GFLOPs
CoBEVT 11.14M 213
V2X-ViT 14.17M 287
Where2comm 8.69M 177
CoAlign 12.94M 96
ParCon(Ours) 5.57M 85
Table 3: Performance comparison between ParCon and ParCon-S at AP@0.7. ParCon-S is the sequential architecture model using the same sub-modules of ParCon.
Model V2XSet OPV2V DAIR-V2X
ParCon 0.707 0.649 0.438
ParCon-S 0.690 0.642 0.430
Figure 4: Robustness in various noise ranges on V2XSet, OPV2V, and DAIR-V2X.
Figure 5: Robustness comparison between ParCon and ParCon-S in various noise ranges on V2XSet. ParCon-S is the sequential architecture model using the same sub-modules of ParCon.
@@ 4.2 Quantitative Evaluation
Detection Accuracy. Table 1 compares the proposed ParCon with the other SOTA methods on the simulation datasets, V2XSet and OPV2V, and the real-world dataset, DAIR-V2X. We also consider single-agent detection (No Fusion) and Late Fusion. We observe the following. i) In the simulated datasets, in the mild noise setting, ParCon achieves the SOTA accuracy increased by 5.83% compared with Where2comm on V2XSet and by 5.52% compared with V2X-ViT on OPV2V at AP@0.7. ii) Even in the real-world dataset, in the mild noise setting, ParCon achieves the SOTA accuracy increased by 6.91% compared with V2X-ViT at AP@0.7. In summary, our proposed collaborative perception model, ParCon, outperforms the SOTA models on all the datasets. iii) Our model achieves the highest performance at AP@0.7, except in the mild noise setting of the DAIR-V2X dataset. In this specific case, the No Fusion approach shows the highest performance. This is because the DAIR-V2X dataset includes only one vehicle and one infrastructure, providing limited opportunity for improvement through collaboration. Thus, all the fusion methods, including ours, do not outperform the No Fusion approach in this scenario.
Model Efficiency. Table 3 compares model efficiency in terms of # Params and GFLOPs. The # Params in ParCon is reduced by 56.95% compared with CoAlign and by 60.69% compared with V2X-ViT. The GFLOPs of ParCon is reduced by 11.46% compared with CoAlign and 70.38% compared with V2X-ViT. The efficiency stems from our CCL, which compress the channel size from 256 to 64.
Noise Robustness. In Figure 4, we compare the noise robustness on the V2XSet, OPV2V, and DAIR-V2X datasets. ParCon always outperforms the SOTA models across all the datasets, even in the real-world dataset DAIR-V2X. At the same time, ParCon shows strong noise robustness (the lowest accuracy drop) among the SOTA models on the simulated dataset. From zero noise to the maximum noise value, the accuracy (AP@0.7) of ParCon decreases by 4.90%/3.08%/10.01% on V2XSet and by 3.78%/4.32%/13.41% on OPV2V under latency/heading noise/localization noise, respectively. In contrast, V2X-ViT experiences larger drops of 9.78%/3.38%/13.62% on V2XSet and 7.07%/3.95%/15.18% on OPV2V. Similarly, Where2comm shows even greater declines, with decreases of 8.42%/5.37%/18.42% on V2XSet and 5.48%/8.12%/24.55% on OPV2V. You can see detailed values in Appendix A.5.
Importance of Parallel Connection. To identify the effectiveness of parallel connection, we design the sequential architecture model, ParCon-S, which uses the same sub-modules of ParCon and connects them in series. Similar to ParCon, ParCon-S compresses the multi-agent feature with CCL and uses a fully connected layer to restore the channel size of the output feature to match the channel size of the original feature. Table 3 and Figure 5 show the comparison in terms of detection accuracy and noise robustness between ParCon and ParCon-S, respectively. As shown in Table 3, the parallel model always outperforms the sequential model at AP@0.7. The parallel model enhances the performance by 2.38%, 1.17%, and 1.82% compared to the sequential model on the V2XSet, OPV2V, and DAIR-V2X datasets, respectively. Furthermore, as shown in Figure 5, the parallel model outperforms the sequential model across various noise levels. In the harsh noise setting, the accuracy of ParCon drops by 4.91%/3.08%/10.08% under latency/heading noise/localization noise, respectively, while ParCon-S experiences larger drops of 5.46%/3.51%/11.01% (See Appendix A.5 for detail values). This comparison demonstrates that the parallel connection consistently outperforms the sequential connection across all datasets and noise settings.
@@ 4.3 Qualitative Evaluation
[Uncaptioned image] Figure 6: Visualization of the weights of CCL.
Table 4: Summation of the absolute value of CCL weights in the perfect setting.
Section Depth 1 Depth 2 Depth 3
Sec.1 190.153 104.393 59.076
Sec.2 190.497 117.089 71.186
Sec.3 191.221 142.863 121.540
Sec.4 189.761 138.970 81.123
Table 5: Summation of the absolute value of CCL weights under the noise setting.
Section Depth 1 Depth 2 Depth 3
Sec.1 228.254 103.984 48.986
Sec.2 226.726 117.295 70.545
Sec.3 224.471 172.468 145.967
Sec.4 221.155 158.333 84.261
Investigation of CCL Weights. At CCL, the concatenated intents ($\mathbf{M}^{\mathbf{AA}}\parallel\mathbf{M}^{\mathbf{SA}}\parallel\mathbf{M}^{\mathbf{SC}}\parallel\mathbf{M^{\prime}}$) are fused each other channel-wise. As we visualize the weights of CCL in Figure 6, it is noteworthy that after the second depth, the weights appear divided into four sections corresponding to the concatenated intents. Also, we observe that i) as shown in Table 5 and Table 5, the summation of the absolute weights in the third section of CCL, corresponding to the intent of S-Conv, shows the highest score, and ii) the summation of the absolute weights in the third section of CCL in the fine-tuned model is higher than in the perfect model. These observations suggest that S-Conv, which generates spatial local information, plays a crucial role in noise rejection, and that CCL leverages this by focusing more on the intent of S-Conv. In essence, CCL aggregates the intents of specific effective modules to enhance overall performance and compensate for the limitations of other modules.
Visualization of Detection Results. Figure 7 shows the detection visualization of CoBEVT, V2V-ViT, and Where2comm in V2V and V2X scenarios of V2XSet. Our model accurately predicts bounding boxes that are well-aligned with the ground truth and successfully detects objects under occlusion. In contrast, other models display misalignments and fail to detect objects when occlusion occurs.
4.4 Ablation Studies
Table 6: Effect of compression ratio at AP@0.7.
Rate Perfect Mild Noise # Params GFLOPs
$\times$1 0.839 0.670 19.69M 451
$\times$2 0.838 0.687 8.74M 161
$\times$4 (Ours) 0.850 0.707 5.57M 85
$\times$8 0.851 0.689 5.00M 64
Table 7: Effect of HRPE in the mild noise setting on V2XSet.
HRPE AP@0.5 AP@0.7
0.850 0.707
0.844 0.694
Table 8: Effect of sub-modules in the mild noise setting.
A-Att S-Att S-Conv Original AP@0.5 AP@0.7
0.850 0.707
0.774 0.645
0.848 0.702
0.813 0.663
0.842 0.683
Table 9: Ablation study of S-Att. Whether to apply the S-Att or vanilla model (ViT) in the mild noise setting on V2XSet.
S-Att AP@0.5 AP@0.7 # Params GFLOPs
0.850 0.707 5.57M 85
0.809 0.660 11.79M 75
Figure 7: Visualization of detected boxes in the V2XSet Dataset. The green boxes represent the ground truth, while the red ones represent prediction. Our model ParCon shows more precise detection.
CCL compression ratio. We train ParCon with various compression rates in CCL to validate whether the reduction in feature size causes information loss. Table 7 shows that ParCon with the rate of $\times$4 yields the best performance at AP@0.7. Compared with Parcon with the rate of $\times$1, Parcon with the rate of $\times$4 yields the accuracy enhanced by 5.50% in the mild noise setting, and the number of parameters and GFLOPs decreased by 71.68% and 81.15%, respectively. Based on this result, we select the rate of $\times$4 as the final setting.
Effect of HRPE. We use HRPE in A-Att to help the module understand the features of the other agents. Table 7 shows that HRPE enhances the performance of the ParCon in the mild noise setting. We notice that HRPE does not contribute much in the perfect setting. By appropriately setting the two hyperparameters for dividing $\theta$ and $d$, HRPE makes our model robust to noise.
Effect of sub-modules. We present the ablation study results in Table 9. The inter-agent complementary information, which is related to A-Att, affects the performance of ParCon significantly. The performance AP@0.7 decreases by 8.76% by eliminating A-Att. The next contributing information is due to S-Conv; if S-Conv is absent, the performance AP@0.7 decreases by 6.12%.
Effect of S-Att. In Table 9, we compare our S-Att module with ViT (Dosovitskiy et al. [2020]), which is vanilla model of S-Att. Although using S-Att reduces the number of parameters by 52.76%, it enhances the performance of AP@0.7 by 7.10% in the mild noise setting.
5 Conclusion
This paper has proposed a new collaborative perception model, ParCon, featuring the parallel connection of sub-modules. With the parallel architecture, ParCon effectively handles communication noises, such as latency, heading noise, and localization noise. Also, ParCon is efficient in terms of the number of parameters and GFLOPs as it uses CCL and S-Att. CCL reduces the feature size to make our fusion module highly lightweight with enhanced accuracy under the noise. We adopt DiNAT (Hassani & Shi [2022]) to capture the spatial global context in S-Att, which enhances the performance while simultaneously reducing the number of parameters compared to the vanilla model, ViT (Dosovitskiy et al. [2020]). Our work provides an important guideline for designing new collaborative perception architectures by demonstrating the benefits of parallel architecture, CCL, and S-Att, as well as their effectiveness and efficiency in handling communication noise.
Limitation and future work. To focus on the effectiveness of the parallel architecture, we use the same sub-modules as in other SOTA models, with the exception of S-Att. In future work, we plan to design advanced sub-modules that can enhance synergy with the parallel architecture.
Appendix A Appendix
A.1 Dataset.
We use the V2XSet Xu et al. [2022b], OPV2V Xu et al. [2022c], and DAIR-V2X Yu et al. [2022] datasets to train and validate models in both V2V and V2X scenarios.
V2XSet. V2XSet is a simulated dataset supporting V2X perception, co-simulated using CARLA Dosovitskiy et al. [2017] and OpenCDA Xu et al. [2021]. It comprises 73 scenes with a minimum of 2 to 5 connected agents and includes 11K 3D annotated LiDAR point cloud frames. The training, validation, and testing sets consist of 6.7K, 2K, and 2.8K frames, respectively.
OPV2V. OPV2V designed for multi-agent V2V perception. Each frame typically comprises approximately 3 CAVs, with a minimum of 2 and a maximum of 7. It includes 10.9K LiDAR point cloud frames with 3D annotations. The training, validation, and testing splits include 6.8K, 2K, and 2.2K frames, respectively.
DAIR-V2X. DAIR-V2X is a real-world dataset for collaborative perception. The dataset includes 9K frames from a vehicle and a road infrastructure, which is equipped with both a LiDAR and a 1920x1080 camera. The LiDAR of infrastructure is 300-channel, and the vehicle’s is 40-channel.
A.2 Comparison Models.
We select CoBEVT Xu et al. [2022a], V2X-ViT Xu et al. [2022b], Where2comm Hu et al. [2022], CoAlign Lu et al. [2023] to compare ParCon.
CoBEVT. CoBEVT presents a framework for multi-agent multi-camera perception that collaboratively produces BEV map predictions using both camera and LiDAR.
V2X-ViT. V2X-viT attempts to utilize heterogeneous agents, such as vehicles and road infrastructure, and solve the problem of heterogeneity between them using a transformer with a graph structure.
Where2comm. Where2comm introduces a spatial confidence map, enabling agents to share only spatially sparse yet perceptually important information.
CoAlign. CoAlign introduces a hybrid framework combining intermediate and late fusion, and leverages an agent-object pose graph to align and optimize relative poses between agents, making the system more robust to noise introduced by pose estimation errors.
A.3 Noise Settings and Trained Models.
We define four different types of noise as shown in Table 11, and three different types of trained model as shown in Table 11. The latency is based on the total time delay used in Xu et al. [2022b] and is set to follow uniform distribution $U(0,t_{\mathrm{lag.}})$ with the maximum latency $t_{\mathrm{lag.}}$. The heading and localization noises follow normal distributions with zero mean and standard deviations $\sigma_{\mathrm{hdg.}}$ and $\sigma_{\mathrm{loc.}}$, that is, $\mathcal{N}(0,\sigma_{\mathrm{hdg.}}^{2})$ and $\mathcal{N}(0,\sigma_{\mathrm{loc.}}^{2})$, respectively. We use three different noise settings: Perfect, Mild Noise, and Harsh Noise. In the perfect setting, we let $t_{\mathrm{lag.}}=0,\sigma_{\mathrm{hdg.}}=0,\sigma_{\mathrm{loc.}}=0$, and in the mild noise setting, we let $t_{\mathrm{lag.}}=200,\sigma_{\mathrm{hdg.}}=0.2,\sigma_{\mathrm{loc.}}=0.2$. In the harsh noise setting, we use different values of $t_{\mathrm{lag.}}\in[0,500],\sigma_{\mathrm{hdg.}}=[0,1.0],\sigma_{\mathrm{loc.}}=[0,0.5]$. Also, we train the three types of models based on the noise setting used in the training: Perfect, Noisy, and Fine-tuned. The perfect model is trained with the perfect setting, the noisy model with the simple noise setting, and the fine-tuned model is fine-tuning the perfect model with the mild noise setting.
Table 10: Various types of noise setting. Noise Setting Latency Hdg. Noise Loc. Noise Perfect 0 0 0 Simple Noise 100 $\mathcal{N}(0,0.2^{2})$ $\mathcal{N}(0,0.2^{2})$ Mild Noise $U(0,200)$ $\mathcal{N}(0,0.2^{2})$ $\mathcal{N}(0,0.2^{2})$ Harsh Noise $U(0,t_{\mathrm{lag.}})$ $\mathcal{N}(0,\sigma_{\mathrm{hdg.}}^{2})$ $\mathcal{N}(0,\sigma_{\mathrm{loc.}}^{2})$ $t_{\mathrm{lag.}}\in[0,500]$ $\sigma_{\mathrm{hdg.}}\in[0.0,1.0]$ $\sigma_{\mathrm{loc.}}\in[0.0,0.5]$ Table 11: Various types of trained model. Training Type Method Perfect Training model in the perfect setting. Noise Training model in the simple noise setting. Fine-tuned First, training model in the perfect setting. Then, fine-tuning model in the mild noise setting.
A.4 Attention Weights Visualization.
In Figure 8, we visualize the attention weights in A-Att, where the brighter colors indicate the areas that the ego agent focuses on. Some observations to notice are as follows: i) Our models show a clear difference between the regions that are paid attention to or not, compared to the V2X-ViT’s attention weights (Figure 8-(b)). ii) Our models have different attention tendencies. For ParCon in Figure 8-(d), the ego agent relies on its own confidence regions where its point clouds are dense and not occluded. Also, it relies on the aux agent’s confidence regions to incorporate wide-range information. In the sequential model in Figure 8-(c), the ego agent focuses on the areas where objects densely exist. The tendencies are more explicitly identified in the infrastructure data, which has a wider detection range but is sparse and can be noisy at far distances. In ParCon, the ego agent takes the infra data relatively less, while the ego agent in the sequential model utilizes the infrastructure data heavily because the infra data has lower confidence than the data from nearby aux agents.
(a) V2X Scenario        (b) V2X-ViT        (c) ParCon-S        (d) ParCon